OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

terenzreignz (terenzreignz):

The point-slope form of the line is given by y - k = m(x - h) Where m is the slope and (h,k) is a point on the line. So we are given a point on the line, (1, -2), so all we need to do is get the slope. It is said that the slope is the same as the slope of the line y - 4 = (2/3)(x - 3) Which incidentally, already happens to be in the point-slope form :D That means 2/3 is the slope. So we have our slope (2/3) and we have our point (1, -2) Plugging them into the point slope form, we get y + 2 = (2/3)(x - 1) The standard form is Ax + By = C y + 2 = (2/3)x - (2/3) (2/3)x + y = -2 - (2/3) (2/3)x + y = -(8/3) Cheers :D